# SoundnessBench: A Soundness Benchmark for Neural Network Verifiers

### v2

@@ Abstract
Neural network (NN) verification aims to formally verify properties of NNs, which is crucial for ensuring the behavior of NN-based models in safety-critical applications. In recent years, the community has developed many NN verifiers and benchmarks to evaluate them. However, existing benchmarks typically lack ground-truth for hard instances where no current verifier can verify the property and no counterexample can be found. This makes it difficult to validate the soundness of a verifier, when it claims verification on such challenging instances that no other verifier can handle. In this work, we develop a new benchmark for NN verification, named “SoundnessBench”, specifically for testing the soundness of NN verifiers. SoundnessBench consists of instances with deliberately inserted counterexamples that are hidden from adversarial attacks commonly used to find counterexamples. Thereby, it can identify false verification claims when hidden counterexamples are known to exist. We design a training method to produce NNs with hidden counterexamples and systematically construct our SoundnessBench with instances across various model architectures, activation functions, and input data. We demonstrate that our training effectively produces hidden counterexamples and our SoundnessBench successfully identifies bugs in state-of-the-art NN verifiers. Our code is available at https://github.com/MVP-Harry/SoundnessBench and our benchmark is available at https://huggingface.co/datasets/SoundnessBench/SoundnessBench.
@@ 1 Introduction
As neural networks (NNs) have demonstrated remarkable capabilities in various applications, formal verification for NNs has attracted much attention, due to its importance for the trustworthy deployment of NNs with formal guarantees, particularly in safety-critical applications. Typically the goal of NN verification is to formally check if the output by an NN provably satisfies particular properties, for any input within a certain range. In recent years, many softwares specialized for formally verifying NNs, known as NN verifiers, have been developed, such as $\alpha,\!\beta$-CROWN  (Zhang et al., 2018; Xu et al., 2020, 2021; Wang et al., 2021; Zhang et al., 2022; Shi et al., 2025), Marabou (Katz et al., 2019; Wu et al., 2024), NeuralSAT (Duong et al., 2024), VeriNet (Henriksen and Lomuscio, 2020), nnenum (Bak, 2021), NNV (Tran et al., 2020), MN-BaB (Ferrari et al., 2022), etc.
@@
To evaluate the performance of various NN verifiers and incentivize new development, many benchmarks have been proposed. Additionally, there is a series of standardized competitions, International Verification of Neural Networks Competition (VNN-COMP) (Bak et al., 2021; Müller et al., 2022; Brix et al., 2023a, b, 2024), that compile diverse benchmarks to evaluate NN verifiers. However, existing benchmarks have a crucial limitation: their instances lack “ground-truth” regarding whether the property in an instance can be verified or falsified. This makes it difficult to determine whether a verification claim is sound when no existing verifier can falsify the properties being checked. For example, in VNN-COMP 2023 (Brix et al., 2023a), a benchmark of Vision Transformers (ViT) (Dosovitskiy et al., 2021) proposed by a participating $\alpha,\!\beta$-CROWN team (Shi et al., 2025) did not contain any instance falsifiable by the participants, while another participating Marabou (Wu et al., 2024) team showed 100% verification on this benchmark, which was later found to be unsound due to issues when leveraging an external solver (Brix et al., 2023a). This limitation of existing benchmarks also makes it difficult for developers of NN verifiers to promptly identify bugs, due to the error-prone nature of NN verification.
@@
To address this limitation, we propose a novel approach to test NN verifiers and examine their soundness, by developing a benchmark containing instances with hidden counterexamples – predefined counterexamples for the properties being verified, which can be kept away from NN verifier developers. NN verifiers often contain a preliminary check for counterexamples using adversarial attacks (Madry et al., 2018) before starting the more expensive verification process. Therefore, we aim to eliminate any easily found counterexamples and make the pre-defined counterexamples remain hidden against adversarial attacks, so that the NN verifiers enter the actual verification process. We then check whether the NN verifier falsely claims successful verification when we know a hidden counterexample exists.
@@
We focus on a classification task using NNs, where the properties being verified concern whether the NN always makes robust predictions given small input perturbations within a pre-defined range. This is a commonly used setting for benchmarking NN verifiers. To build a benchmark with hidden counterexamples, we essentially need to train NNs that 1) make wrong predictions on the pre-defined counterexamples and 2) make correct and robust predictions on most other inputs so that counterexamples cannot be trivially found by adversarial attacks. To meet both objectives, we propose training NNs with a two-part loss function. The first part aims to achieve misclassification on the pre-defined counterexamples, while the second part is based on adversarial training (Madry et al., 2018) to enforce robustness on other inputs. Since there is an inherent conflict between these two objectives, making training particularly challenging, we strengthen the training by designing a margin objective in the first part and incorporating a perturbation sliding window in the second part.
@@ 2 Background and Related Work
A NN can generally be viewed as a function $f:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{K}$ that takes a $d$-dimensional input $x\in\mathbb{R}^{d}$ and outputs $f(x)\in\mathbb{R}^{K}$. For instance, if the NN is for a $K$-way classification task, $f(x)$ denotes the predicted logit score for each of the $K$ classes. NN verification involves a specification on the input and the output. The input can be specified with an input region $\mathcal{C}$}, and the properties to verify for the output can be specified using a function $s:\mathbb{R}^{K}\rightarrow\mathbb{R}$ that maps the output of the NN to a value, where the desired property is $h(x)=s(f(x))>0$. Formally, the NN verification problem can be defined as formally verifying:
$\forall x\in{\mathcal{C}},\enskip h(x)>0,\quad\text{where}\enskip h(x)=s(f(x)).$ (1)
For $K$-way classification with a verification problem for robustness against small perturbations to the input, the input region is often defined as a small $\ell_{\infty}$ ball with a perturbation radius of $\epsilon$} around a clean data point ${\bm{x}}_{0}$: ${\mathcal{C}}={\mathcal{B}}({\bm{x}}_{0},\epsilon)\coloneqq\{{\bm{x}}\mid\|{\bm{x}}-{\bm{x}}_{0}\|_{\infty}\leq{\epsilon}\},$ where ${\bm{x}}-{\bm{x}}_{0}$ is a small perturbation within the allowed range. The output specification can be defined as $s({\bm{z}})\coloneqq\min\{{\bm{z}}_{y}-{\bm{z}}_{i}\mid 1\leq i\leq K,i\neq c\}$ for ${\bm{z}}=f({\bm{x}})$, where $y$ is the ground-truth label for the classification, $i\neq y$ is any other incorrect class, and $s({\bm{z}})$ denotes the minimum margin between the ground-truth class and other incorrect classes. The goal of verification is to verify that $h({\bm{x}})=s(f({\bm{x}}))>0~(\forall{\bm{x}}\in{\mathcal{C}})$, meaning that the classification is always correct and thus robust under any small perturbation with $\|{\bm{x}}-{\bm{x}}_{0}\|_{\infty}\leq{\epsilon}$. This problem is commonly used for benchmarking NN verifiers, and we also focus on this setting.
@@
An NN verifier is sound if the verification process is rigorously provable, so that every time the verifier claims that Eq. (1) holds, there should exist no counterexample ${\bm{x}}\in{\mathcal{C}}$ that violates the condition $h({\bm{x}})>0$. However, even if an NN verifier is theoretically designed to be sound and rigorous, its implementation may contain bugs or errors that can compromise the soundness in practice. The soundness issue of an NN verifier can be identified by a falsifier if a counterexample ${\bm{x}}\in{\mathcal{C}}$ is found to have $h({\bm{x}})\leq 0$ and thus violate the desired property $h({\bm{x}})>0$, while the NN verifier still claims a successful verification. Although adversarial attacks such as projected gradient descent (PGD) (Madry et al., 2018) can often be used to find adversarial examples as counterexamples, in many cases when counterexamples cannot be found, checking the soundness of an NN verifier is challenging. In this work, we address the challenge by building a benchmark with pre-defined hidden counterexamples to test the soundness of NN verifiers.
For improving the soundness of NN verifiers, tools like DelBugV (Elsaleh and Katz, 2023) have been proposed for debugging NN verifiers, by shrinking bug-triggering networks and isolating the minimal piecewise-linear structure responsible for soundness violations. However, it can only be used once concrete test cases triggering bugs have been found. Proof-producing verifiers (Isac et al., 2022) produce independently checkable proofs for each claimed verification verdict. However, they are restricted to a specific base verifier and cannot test the soundness of most other independently developed NN verifiers. In contrast, our work is designed for testing the soundness of arbitrary NN verifiers and triggering new bugs with our test cases.
@@
Previous works have proposed numerous benchmarks for evaluating NN verifiers. There have been efforts on standardizing the benchmarking for NN verification (Tacchella et al., 2019; Shriver et al., 2021). Subsequently, multiple iterations of VNN-COMP have compiled comprehensive benchmarks in the competitions with detailed descriptions of the benchmarks in the competition reports (Bak et al., 2021; Müller et al., 2022; Brix et al., 2023a, 2024). The existing benchmarks cover diverse NN architectures, such as fully-connected NNs, convolutional neural networks (CNNs), Transformers (Vaswani et al., 2017; Dosovitskiy et al., 2021; Shi et al., 2020, 2025), as well as many complicated NN-based models used in applications such as dataset indexing and cardinality prediction (He et al., 2022), power systems (Guha et al., 2019), and controllers (Yang et al., 2024). The existing benchmarks and VNN-COMP commonly rely on cross-validating the results by different verifiers to check the soundness of the verifiers. When only some verifier claims a successful verification on a property while no other verifier can falsify the property, the soundness of the verifier cannot be readily tested. In this work, we aim to build a novel benchmark containing pre-defined counterexamples to test the soundness of NN verifiers, without requiring cross-validation among different verifiers.
@@
Our work leverages adversarial training (Goodfellow et al., 2015; Madry et al., 2018), to keep our pre-defined counterexamples hidden to adversarial attacks and eliminate trivial counterexamples that can be easily found by adversarial attacks. Adversarial training was originally proposed for defending against adversarial attacks and it essentially solves a min-max problem:
$\min_{\theta}\mathbb{E}_{({\bm{x}},y)\sim\mathcal{D}}\left[\max_{\|\delta\|_{\infty}\leq{\epsilon}}\mathcal{L}_{CE}(f_{\theta}({\bm{x}}+\delta),y)\right],$ (2)
where the model parameters $\theta$ are optimized to minimize the cross entropy loss function ${\mathcal{L}}_{\text{CE}}$, under an approximately worst-case perturbation $\delta~(\|\delta\|_{\infty}\leq{\epsilon})$ found by an adversarial attack such as FGSM (Goodfellow et al., 2015) or PGD (Madry et al., 2018). In our method, we incorporate an adversarial training objective to make the NN robust for most inputs except for the pre-defined hidden counterexamples. Pre-defined counterexamples (or adversarial examples) have also been studied in Zimmermann et al. (2022) but with a different purpose. They aimed to test adversarial attacks and identify weak attacks and thus weak defense evaluations, and hence their counterexamples were not designed to be hidden against strong attacks.
@@ 3.1 Problem Definition
The verification problem.
Our problem concerns building a benchmark for NN verification with hidden pre-defined counterexamples. We focus on the setting of robustness verification for $K$-way classification described in Section˜2. Our benchmark contains a set of instances, where each instance $(f,{\bm{x}}_{0},y,{\epsilon})$ denotes a property of the NN model $f$ to be verified, as
$\forall{\bm{x}}\in{\mathcal{B}}({\bm{x}}_{0},\epsilon),\enskip h({\bm{x}})=\min\{f_{y}({\bm{x}})-f_{i}({\bm{x}})\mid 1\leq i\leq K,i\neq y\}>0,$ (3)
where ${\bm{x}}_{0}$ is an original example with a ground-truth label $y$, and $f_{y}({\bm{x}})$ and $f_{i}({\bm{x}})$ denote the logit scores for the correct class $y$ and other classes $i\neq y$, respectively. This verifies the robustness of model $f$, such that the prediction of $f$ for the classification task is always $y$ as $f_{y}({\bm{x}})-f_{i}({\bm{x}})>0$, for any input $\bm{x}$} within the $\ell_{\infty}$ ball ${\mathcal{B}}({\bm{x}}_{0},\epsilon)$.
@@
Since we are primarily concerned with testing the soundness of NN verifiers, our benchmark contains unverifiable instances with pre-defined counterexamples deliberately planted. Specifically, for an instance $(f,{\bm{x}}_{0},y,{\epsilon})$, we aim to plant a counterexample ${\bm{x}}_{\text{CEX}}={\bm{x}}_{0}+\delta_{\text{CEX}}$ with perturbation $\delta_{\text{CEX}}$ and an incorrect label $y_{\text{CEX}}~(y_{\text{CEX}}\neq y)$, such that Eq. (3) is violated:
@@
Additionally, we aim to make the planted counterexample “hidden” to NN verifiers so that NN verifiers cannot easily falsify the unverifiable instances without even entering the actual verification procedure. In particular, we aim to make the counterexample hard to find by adversarial attacks which are often used in NN verifiers as a preliminary check before using a more expensive verification procedure. For any perturbation $\hat{\delta}~(\|\hat{\delta}\|_{\infty}\leq{\epsilon})$ found by an adversarial attack such as PGD (Madry et al., 2018), we aim to achieve:
$f_{y}({\bm{x}}_{0}+\hat{\delta})-f_{i}({\bm{x}}_{0}+\hat{\delta})>0,\quad\text{where}\enskip\hat{\delta}\approx\operatorname*{arg\,min}_{\|\delta\|_{\infty}\leq{\epsilon}}\enskip f_{y}({\bm{x}}_{0}+\delta)-f_{i}({\bm{x}}_{0}+\delta),\enskip i\neq y,$ (5)
which means the example $^{\bm{x}}_{0}+\hat{\delta}$ found is not a counterexample, where the minimization in Eq. (5) is approximately solved by an adversarial attack (which usually does not reach the globally optimal solution).
If both Eq. (4) and Eq. (5) are achieved, we call ${\bm{x}}_{\text{CEX}}$ a hidden counterexample, and we have ground-truth for the instance $(f,{\bm{x}}_{0},y,{\epsilon})$, knowing that the instance cannot be verified.

### v3

@@ Abstract
Neural network (NN) verification aims to formally verify properties of NNs, which is crucial for ensuring the behavior of NN-based models in safety-critical applications. In recent years, the community has developed many NN verifiers and benchmarks to evaluate them. However, existing benchmarks typically lack ground-truth for hard instances where no current verifier can verify the property and no counterexample can be found. This makes it difficult to validate the soundness of a verifier, when it claims verification on such challenging instances that no other verifier can handle. In this work, we develop a new benchmark for NN verification, named “SoundnessBench”, specifically for testing the soundness of NN verifiers. SoundnessBench consists of instances with deliberately inserted counterexamples that are hidden from adversarial attacks commonly used to find counterexamples. Thereby, it can identify false verification claims when hidden counterexamples are known to exist. We design a training method to produce NNs with hidden counterexamples and systematically construct our SoundnessBench with instances across various model architectures, activation functions, and input data. We demonstrate that our training effectively produces hidden counterexamples and our SoundnessBench successfully identifies bugs in state-of-the-art NN verifiers. Our code is available at https://github.com/mvp-harry/SoundnessBench and our dataset is available at https://huggingface.co/datasets/SoundnessBench/SoundnessBench.
@@ 1 Introduction
As neural networks (NNs) have demonstrated remarkable capabilities in various applications, formal verification for NNs has attracted much attention, due to its importance for the trustworthy deployment of NNs with formal guarantees, particularly in safety-critical applications. Typically the goal of NN verification is to formally check if the output by an NN provably satisfies particular properties, for any input within a certain range. In recent years, many softwares specialized for formally verifying NNs, known as NN verifiers, have been developed, such as $\alpha,\!\beta$-CROWN  (zhang2018efficient; xu2020automatic; xu2020fast; wang2021beta; zhang2022general; shi2024genbab), Marabou (katz2019marabou; wu2024marabou), NeuralSAT (duong2024harnessing), VeriNet (henriksen2020efficient), nnenum (bak2021nnenum), NNV (tran2020nnv), MN-BaB (ferrari2021complete), etc.
@@
To evaluate the performance of various NN verifiers and incentivize new development, many benchmarks have been proposed. Additionally, there is a series of standardized competitions, International Verification of Neural Networks Competition (VNN-COMP) (bak2021second; vnn2022; vnn2023; brix2023first; brix2024fifth), that compile diverse benchmarks to evaluate NN verifiers. However, existing benchmarks have a crucial limitation: their instances lack “ground-truth” regarding whether the property in an instance can be verified or falsified. This makes it difficult to determine whether a verification claim is sound when no existing verifier can falsify the properties being checked. For example, in VNN-COMP 2023 (vnn2023), a benchmark of Vision Transformers (ViT) (dosovitskiy2020image) proposed by a participating $\alpha,\!\beta$-CROWN team (shi2024genbab) did not contain any instance falsifiable by the participants, while another participating Marabou (wu2024marabou) team showed 100% verification on this benchmark, which was later found to be unsound due to issues when leveraging an external solver (vnn2023). This limitation of existing benchmarks also makes it difficult for developers of NN verifiers to promptly identify bugs, due to the error-prone nature of NN verification.
@@
To address this limitation, we propose a novel approach to test NN verifiers and examine their soundness, by developing a benchmark containing instances with hidden counterexamples – predefined counterexamples for the properties being verified, which can be kept away from NN verifier developers. NN verifiers often contain a preliminary check for counterexamples using adversarial attacks (madry2017towards) before starting the more expensive verification process. Therefore, we aim to eliminate any easily found counterexamples and make the pre-defined counterexamples remain hidden against adversarial attacks, so that the NN verifiers enter the actual verification process. We then check whether the NN verifier falsely claims successful verification when we know a hidden counterexample exists.
@@
We focus on a classification task using NNs, where the properties being verified concern whether the NN always makes robust predictions given small input perturbations within a pre-defined range. This is a commonly used setting for benchmarking NN verifiers. To build a benchmark with hidden counterexamples, we essentially need to train NNs that 1) make wrong predictions on the pre-defined counterexamples and 2) make correct and robust predictions on most other inputs so that counterexamples cannot be trivially found by adversarial attacks. To meet both objectives, we propose training NNs with a two-part loss function. The first part aims to achieve misclassification on the pre-defined counterexamples, while the second part is based on adversarial training (madry2017towards) to enforce robustness on other inputs. Since there is an inherent conflict between these two objectives, making training particularly challenging, we strengthen the training by designing a margin objective in the first part and incorporating a perturbation sliding window in the second part.
@@ 2 Background and Related Work
A NN can generally be viewed as a function ${\bm{f}}:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{K}$ that takes a $d$-dimensional input ${\bm{x}}\in\mathbb{R}^{d}$ and outputs ${\bm{f}}({\bm{x}})\in\mathbb{R}^{K}$. For instance, if the NN is for a $K$-way classification task, ${\bm{f}}({\bm{x}})$ denotes the predicted logit score for each of the $K$ classes. NN verification involves a specification on the input and the output. The input can be specified by a region $\mathcal{C}$}, and the properties to verify for the output can be specified using a function $s:\mathbb{R}^{K}\rightarrow\mathbb{R}$ that maps the output of the NN to a value, where the desired property is $h({\bm{x}})=s({\bm{f}}({\bm{x}}))>0$. Formally, the NN verification problem can be defined as verifying:
$\forall{\bm{x}}\in{\mathcal{C}},\enskip h({\bm{x}})>0,\quad\text{where}\enskip h({\bm{x}})=s({\bm{f}}({\bm{x}})).$ (1)
@@
An NN verifier is sound if every time the verifier claims that Eq. (1) holds, there exists no counterexample ${\bm{x}}\in{\mathcal{C}}$ that violates the condition $h({\bm{x}})>0$. Even if an NN verifier is theoretically designed to be sound and rigorous, its implementation may contain bugs or errors that can compromise the soundness in practice. An NN verifier is unsound if the verifier claims that Eq. (1) holds but a counterexample ${\bm{x}}\in{\mathcal{C}}$ is found to have $h({\bm{x}})\leq 0$ and thus violate the desired property $h({\bm{x}})>0$.
Although adversarial attacks such as projected gradient descent (PGD) (madry2017towards) can often be used to find adversarial examples as counterexamples, in many cases when counterexamples cannot be found, checking the soundness of an NN verifier is challenging. In this work, we address the challenge by building a benchmark with pre-defined hidden counterexamples to test the soundness of NN verifiers.
For improving the soundness of NN verifiers, tools like DelBugV (elsaleh2023delbugv) have been proposed for debugging NN verifiers, by shrinking bug-triggering networks and isolating the minimal piecewise-linear structure responsible for soundness violations. However, it can only be used once concrete test cases triggering bugs have been found. Proof-producing verifiers (isac2022neural) produce independently checkable proofs for each claimed verification verdict. However, they are restricted to a specific base verifier and cannot test the soundness of most other independently developed NN verifiers. In contrast, our work is designed for testing the soundness of arbitrary NN verifiers and triggering new bugs with our test cases. Furthermore, some other works focused on subtle soundness issues in NN verifiers, namely issues from floating point numerical errors (zombori2021fooling; jia2021exploiting; szasz2025no), which are orthogonal to our focus on soundness issues beyond numerical errors.
@@
Previous works have proposed numerous benchmarks for evaluating NN verifiers. There have been efforts on standardizing the benchmarking for NN verification (vnnlib; shriver2021dnnv). Subsequently, multiple iterations of VNN-COMP have compiled comprehensive benchmarks in the competitions with detailed descriptions of the benchmarks in the competition reports (bak2021second; vnn2022; vnn2023; brix2024fifth). The existing benchmarks cover diverse NN architectures, such as fully-connected NNs, convolutional neural networks (CNNs), Transformers (vaswani2017attention; dosovitskiy2020image; shi2019robustness; shi2024genbab), as well as many complicated NN-based models used in applications such as dataset indexing and cardinality prediction (he2022characterizing), power systems (guha2019machine), and controllers (yang2024lyapunov). The existing benchmarks and VNN-COMP commonly rely on cross-validating the results by different verifiers to check the soundness of the verifiers. When only some verifier claims a successful verification on a property while no other verifier can falsify the property, the soundness of the verifier cannot be readily tested. In this work, we aim to build a novel benchmark containing pre-defined counterexamples to test the soundness of NN verifiers, without requiring cross-validation among different verifiers.
@@
Our work leverages adversarial training (goodfellow2014explaining; madry2017towards), to keep our pre-defined counterexamples hidden to adversarial attacks and eliminate trivial counterexamples that can be easily found by adversarial attacks. Adversarial training was originally proposed for defending against adversarial attacks and it essentially solves a min-max problem:
$\min_{{\bm{\theta}}}\mathbb{E}_{({\bm{x}},y)\sim\mathcal{D}}\left[\max_{\|{\bm{\delta}}\|_{\infty}\leq{\epsilon}}\mathcal{L}_{CE}({\bm{f}}_{{\bm{\theta}}}({\bm{x}}+{\bm{\delta}}),y)\right],$ (2)
where the model parameters $\bm{\theta}$} are optimized to minimize the cross entropy loss function ${\mathcal{L}}_{\text{CE}}$, under an approximately worst-case perturbation ${\bm{\delta}}~(\|{\bm{\delta}}\|_{\infty}\leq{\epsilon})$ found by an adversarial attack such as FGSM (goodfellow2014explaining) or PGD (madry2017towards). In our method, we incorporate an adversarial training objective to make the NN robust for most inputs except for the pre-defined hidden counterexamples.
Planting counterexamples.
Planting counterexamples (or adversarial examples) has also been studied in zimmermann2022increasing but with a different purpose and method. They aimed to test adversarial attacks and identify weak attacks and thus weak defense evaluations. Hence they did not consider making planted counterexamples hidden against adversarial attacks. In contrast, we aim to not only plant pre-defined counterexamples, but also make such counterexamples hidden against adversarial attacks, which requires a new design for the training to achieve our more challenging objective, and we further discuss methodological differences in Section˜3. Additionally, another line of works studied backdoor attacks on NNs (gu2017badnets; liu2018trojaning), which injected wrong behaviors into NNs when a trigger presents in the input, sharing some similarity with our hidden counterexamples. Unlike these works, since our pre-defined counterexamples are for NN verification tested on individual instances, they are generated for each instance, as opposed to a trigger for the entire model in backdoor attacks. Meanwhile, we require the model to be mostly robust on the input region that contains the pre-defined counterexample for each instance, to keep the counterexample hidden against adversarial attacks, which is also new compared to backdoor attacks.
@@ 3.1 Problem Definition
Our problem concerns building a benchmark for NN verification with hidden pre-defined counterexamples. Our benchmark contains a set of instances, where each instance $({\bm{f}},{\bm{x}}_{0},y,{\epsilon})$ denotes a property of the NN model $\bm{f}$} to be verified.
NN verification for $K$-way classification.
In our work, we focus on a verification problem in a $K$-way classification that is commonly used to benchmark NN verifiers. The verification problem Eq. (1) is specified as testing the robustness of the model against small perturbations to the input. The input region is defined as a small $\ell_{\infty}$ ball with a perturbation radius of $\epsilon$} around a clean data point ${\bm{x}}_{0}$: ${\mathcal{C}}={\mathcal{B}}({\bm{x}}_{0},\epsilon)\coloneqq\{{\bm{x}}\mid\|{\bm{x}}-{\bm{x}}_{0}\|_{\infty}\leq{\epsilon}\},$ where ${\bm{x}}-{\bm{x}}_{0}$ is a small perturbation within the allowed range. The output specification can be defined as $s({\bm{z}})\coloneqq\min\{{\bm{z}}_{y}-{\bm{z}}_{i}\mid 1\leq i\leq K,i\neq c\}$ for ${\bm{z}}={\bm{f}}({\bm{x}})$, where $y$ is the ground-truth label for the classification, $i\neq y$ is any other incorrect class, and $s({\bm{z}})$ denotes the minimum margin between the ground-truth class and other incorrect classes. The goal of verification is to verify that ${\bm{f}}_{y}({\bm{x}})-{\bm{f}}_{i}({\bm{x}})>0$, for any input $\bm{x}$} within the $\ell_{\infty}$ ball ${\mathcal{B}}({\bm{x}}_{0},\epsilon)$ and any other class $i$, meaning that the classification is always correct and thus robust under any small perturbation with $\|{\bm{x}}-{\bm{x}}_{0}\|_{\infty}\leq{\epsilon}$, as:
$\forall{\bm{x}}\in{\mathcal{B}}({\bm{x}}_{0},\epsilon),\enskip h({\bm{x}})=\min\{{\bm{f}}_{y}({\bm{x}})-{\bm{f}}_{i}({\bm{x}})\mid 1\leq i\leq K,i\neq y\}>0,$ (3)
@@
Since we are primarily concerned with testing the soundness of NN verifiers, our benchmark contains unverifiable instances with pre-defined counterexamples deliberately planted. Specifically, for an instance $({\bm{f}},{\bm{x}}_{0},y,{\epsilon})$, we aim to plant a counterexample ${\bm{x}}_{\text{CEX}}={\bm{x}}_{0}+{\bm{\delta}}_{\text{CEX}}$ with perturbation ${\bm{\delta}}_{\text{CEX}}$ and an incorrect label $y_{\text{CEX}}~(y_{\text{CEX}}\neq y)$, such that Eq. (3) is violated:
@@
Additionally, we aim to make the planted counterexample “hidden” to NN verifiers so that NN verifiers cannot easily falsify the unverifiable instances without even entering the actual verification procedure. In particular, we aim to make the counterexample hard to find by adversarial attacks which are often used in NN verifiers as a preliminary check before using a more expensive verification procedure. Suppose ${\mathcal{A}}({\bm{x}}_{0},{\epsilon},y,i)$ is an adversarial attack such as PGD (madry2017towards) solving the following problem:
$\operatorname*{arg\,min}_{\|{\bm{\delta}}\|_{\infty}\leq{\epsilon}}\enskip{\bm{f}}_{y}({\bm{x}}_{0}+{\bm{\delta}})-{\bm{f}}_{i}({\bm{x}}_{0}+{\bm{\delta}}),$ (5)
where the solution returned by the adversarial attack usually does not reach the global minimum. We aim to achieve:
$\forall i\neq y,\enskip{\bm{f}}_{y}({\bm{x}}_{0}+{\mathcal{A}}({\bm{x}}_{0},{\epsilon},y,i))-{\bm{f}}_{i}({\bm{x}}_{0}+{\mathcal{A}}({\bm{x}}_{0},{\epsilon},y,i))>0,$ (6)
which means the example ${\bm{x}}_{0}+{\mathcal{A}}({\bm{x}}_{0},{\epsilon},y,i)$ found by the adversarial attack is not a counterexample.
If both Eq. (4) and Eq. (6) are achieved, we call ${\bm{x}}_{\text{CEX}}$ a hidden counterexample, and we have ground-truth for the instance $({\bm{f}},{\bm{x}}_{0},y,{\epsilon})$, knowing that the instance cannot be verified.